OpenStudy (anonymous):

A 15.0 kg block is pulled by a horizontal force of 30.0 N along a rough horizontal surface at constant velocity. What is the coefficient of friction?

OpenStudy (anonymous):



OpenStudy (anonymous):

What is muk?

OpenStudy (anonymous):

\[30 - \mu _{k}*N =0\] \[mg=N\] \[30 - \mu _{k}*mg =0\]

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

ooh ok.. thats better.

OpenStudy (anonymous):

So 30 - muk * 15.0=0 right since mg is equals to newton or nope? @Algebraic!

OpenStudy (anonymous):

yep, you're right so far...

OpenStudy (anonymous):

you forgot 'g' actually...

OpenStudy (anonymous):

m*g

OpenStudy (anonymous):

you just have mass... need a '9.8 m/s^2 ' in there too.

OpenStudy (anonymous):



OpenStudy (anonymous):

do wee need to solve the gravity or nope? because its saying "pulled by a horizontal force" Do you know what i mean right?

OpenStudy (anonymous):

solve: \[30 - \mu _{k}*15*9.8 = 0\] for mu k

OpenStudy (anonymous):

muk= -4425? @Algebraic!

OpenStudy (anonymous):

30/(15*9.8)

OpenStudy (anonymous):

use your algebra skills. this is what you trained for.

OpenStudy (anonymous):

Sorry. xD I got 0.20 My teacher gave us the gravity=9.81

OpenStudy (anonymous):

yes .204 is right.

OpenStudy (goformit100):

:)

OpenStudy (anonymous):

just making sure: the formula is \[F=\mu*N \]?? and then why you put equals zero at the end? i just want to know?

OpenStudy (anonymous):

the 'formula' is F=ma which means net force = ma which means add up all the forces and the sum is ma here there's two forces, the applied force and the friction force and they add up to the acceleration (which is zero) times the mass... Force from friction is muk*Normal since it's a 'flat' surface Normal = mg so: applied force - friction force = 15*0 30 - muk*mg =0

OpenStudy (anonymous):

Thank you. So when its asking for flat surface or what ever is that called it will be Normal = mg the Mass times the gravity right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure!